I. You are developing an "X-bar" chart based on sample means. You know the standard deviation of the sample means is 4, the desired confidence level is 99 percent, and the average of the sample means is 20. Which of the following is your LCL? (Note that Z-value will be 2.576)

A. 36
B. 24
C. 9.68
D. 16.79
E. 30.32

II. You want to develop a three-sigma "R" chart. You know the average range is 12 based on several samples of size 6. Which of the following is the resulting LCL? For n=6, D3=0 and D4=2.004
A. 20.0
B. 18.3
C. 7.02
D. 5.6
E. 0.0

III. Using the cut-and-try method for aggregate operations planning we can determine the production requirement in units of product. If the beginning inventory is 100 units, the demand forecast is 1,200, and the necessary safety stock is 20% of the demand forecast, which of the following is the production requirement?
A. 1,200
B. 1,300
C. 1,340
D. 1,500
E. 1,540

IV. Using the cut-and-try method for aggregate operations planning we can determine the production requirement in units of product. If the beginning inventory is 500 units, the demand forecast is 1,000, and the necessary safety stock is 10% of the demand forecast, which of the following is the production requirement?
A. 1,000
B. 600
C. 550
D. 450
E. 100

V. Which of the following is the planned-order release using an MRP program if the gross requirement is 5,000, inventory on hand is 1,200 and scheduled receipts are 800?
A. 7,000
B. 4,200
C. 3,000
D. 2,000
E. 1,200
Answer #1

I)

Given

= 20

Standard deviation = 4

confidence level = 99 percent

Z = 2.576

LCL = - Z* = 20 - 4*2.58 = 9.68

C. 9.68

II)

LCL = D3 *

LCL = 0*12 = 0

E. 0.0

III)

Beginning inventory = 100 units

Demand forecast = 1,200

Safety stock = 20% of the demand forecast = 0.2*1200 = 240

Production requirement = Demand Forecast + Safety Stock – Beginning Inventory = 1200 + 240 - 100 = 1340

C. 1,340

IV)

Beginning inventory = 500 units

Demand forecast = 1,000

Safety stock = 10% of the demand forecast = 0.10*1000 = 100

Production requirement = Demand Forecast + Safety Stock – Beginning Inventory = 1000 + 100 - 500 = 600

B. 600